OpenStudy (yrivers36):

Find the radian measure and complement (in radians)of an angle measuring 65 degrees.

OpenStudy (anonymous):

mutliply 65 degrees by pi/180 to get radians

OpenStudy (anonymous):

\[65^{o}*\pi/180=\frac{13\pi}{36} \ radians|\]

OpenStudy (yrivers36):

I got that thank you. Now how do I find the complement

OpenStudy (anonymous):

the complement of 65 degrees is 90-65 = 35 degrees, or in radians\[\frac{18 \pi}{36}-\frac{13 \pi}{36}=\frac{5 \pi}{36}\]

OpenStudy (anonymous):

2*pi radians = 1 circle 360 degree = 1 circle 2*pi radians = 360 degrees pi radians = 180 degrees 1 radian = 180/pi degrees so build the conversion factor based on this equivalence statement

OpenStudy (yrivers36):

Mandolino how did you get that last problem

OpenStudy (anonymous):

pi/2 is the radian measure for 90 degrees; i just converted it to a denominator of 36 so i could subtract the 13pi/36